Given two specific terms of geometric sequence, find the common ratio and first term. Show your solution
a2=4 ; a5=32

Answer by greenestamps(13334)



a(5)=a(2)*r^3

32=4r^3
r^3=8
r=2

a(1)=a(2)/r = 4/2 = 2

ANSWERS:
common ratio r=2
1st term a(1)=2